# A Review on Knowledge Graphs for Healthcare: Resources, Applications, and Promises

### v4

@@ 4.3. Clinical Decision Support
However, the sparsity of EHR data typically allows for only a small fraction of medical codes to be learned effectively, thereby restricting the ability of deep learning approaches. To overcome this drawback, knowledge graphs have been applied to incorporate prior medical knowledge for these deep learning models, which augment the representation of medical codes to better support the downstream prediction tasks.
4.3.1. Intermediate Steps to Advance Prediction Models
@@
ICD Coding aims to extract diagnosis and procedure codes from clinical notes, which often consisted of raw text (Mullenbach et al., 2018; Zhang et al., 2020; Vu et al., 2021; Dong et al., 2022). It is often challenging, as the size of the candidate target codes can be large and the distribution of the codes is often long-tailed (Kim and Ganapathi, 2021). To overcome this, Xie et al. (Xie et al., 2019) and Cao et al. (Cao et al., 2020a) propose to leverage knowledge graphs as distant supervision (Min et al., 2013; Liang et al., 2020), and inject the label information via structured knowledge graph propagation by leveraging graph convolution networks (Kipf and Welling, 2016) to learn the correlations among medical codes. Besides, Lu et al. (Lu et al., 2020) propose to leverage knowledge graphs and the co-occurrence graph among clinical nodes simultaneously with a knowledge aggregation module to boost the ICD coding performance. Ren et al. (Ren et al., 2022) design a learning curriculum based on the hierarchical structure of the code to address the highly imbalanced label distribution issue and balance between frequent and rare labels. Overall, injecting additional knowledge with graph neural networks offers a way to mitigate the imbalanced label distribution issue and thus better.
@@
Entity and Relation Extraction from Health Records. Health records contain rich unstructured or semi-structured data, making it difficult for clinicians to access and analyze relevant information. Entity and relation extraction helps convert this unstructured text into structured data that can be more easily processed, understood, and utilized. Specifically, entity extraction aims to identify entity mentions from clinical-free texts. There are two key steps for entity extraction, i.e., named entity recognition (NER) and disambiguation (NED). By leveraging additional knowledge graphs, Varma et al. (Varma et al., 2021) transfer structural knowledge from the knowledge base to the medical domain, improving rare entities’ disambiguation accuracy. Yuan et al. (Yuan et al., 2023) inject additional knowledge from the knowledge graphs for entity linking and proposed two additional strategies, namely Post-pruning and Thresholding, to improve the efficiency and remove the effect of unlinkable entity mention. Fries et al. (Fries et al., 2021) leverage clinical ontologies to provide weak supervision sources to create additional training data for clinical entity disambiguation. Besides, relation extraction aims to identify and classify relationships between entities in unstructured text, which facilitates understanding complex biological processes, drug interactions, and disease mechanisms. To incorporate the external knowledge graph, several works (Fei et al., 2021; Roy and Pan, 2021) proposed additional post-training steps to align the language models with biomedical knowledge. Hong et al. (Hong et al., 2021) construct embeddings for a wide range of codified concepts from EHRs to identify relevant features related to a disease of interest, and Lin et al. (Lin et al., 2022) design a co-training scheme to jointly learn from text and knowledge graphs for extracting and classifying disease-disease relations. In summary, fusing knowledge graphs with language models can flexibly accommodate missing data types and bring additional performance gains, especially for those rare entities and relations.
4.3.2. Evidence Generation for Risk Prediction Models
@@
Disease Prediction aims to predict the potential diseases of a given patient with his past clinical records. To assist the diagnosis with additional knowledge, GRAM (Choi et al., 2017) and KAME (Ma et al., 2018) utilize a medical ontology (Dubberke et al., 2006) where the leaf nodes are the medical codes found in EHR data, and their ancestors are more general categories. By incorporating information from medical ontologies into deep learning models via neural attention, these approaches learn better embeddings for different medical concepts to alleviate the data scarcity bottleneck. (Yin et al., 2019; Zhang et al., 2019) further consider the domain-specific knowledge graph KnowLife (Ernst et al., 2015) to enrich the embeddings of medical entities with their neighbors on the knowledge graph. These approaches mainly directly update the embeddings of different concepts to improve the feature learning, but may be at the risk of ignoring the high-level order information from the knowledge graph. To tackle this drawback, Ye et al. (Ye et al., 2021a) explicitly exploit paths in KG from the observed symptoms to the target disease to model the personalized information for diverse patients with a relational-guided attention mechanism. Xu et al. (Xu et al., 2021) design a self-supervised learning approach to pre-train a graph attention network for learning the embedding of medical concepts and completing the knowledge graph simultaneously. These approaches better harness the structure information, and often lead to better performance than the pure embedding-based knowledge integration techniques.
@@
Treatment Recommendation aims to recommend personalized medications to patients based on their individual health conditions, which can help physicians select the most effective medications for their patients, and improve treatment outcomes (Zhang et al., 2017; Bhoi et al., 2021; Shang et al., 2019b). To effectively exploit external knowledge, Shang et al. (Shang et al., 2019a) use drug ontologies to design additional pretraining loss and directly improve the representation of drugs, and several studies (Wu et al., 2022a; Tan et al., 2022) attempt to extract the additional drug interaction graphs to model the negative side effects of specific drug pairs and reduce the possibility of recommending negative drug-drug interaction combinations. Besides, Wu et al. (Wu et al., 2022b) leveraged ontologies to improve the drug representations, and facilitates drug recommendation under a more challenging few-shot setting.
4.4. Public Health
@@
Public Health research can significantly benefit from HKGs. Knowledge graphs can help organize, structure, and formalize extensive information from diverse and heterogeneous sources. This allows researchers to analyze data, reason about factors, and make decisions on a larger scale.
@@
Epidemiology. The field of epidemiology has seen an increase in the use of knowledge graphs to analyze and understand the spread of diseases. A study conducted by Gao et al. (Gao et al., 2023) analyzes the research hotspots and development trends of wastewater-based epidemiology (WBE) using knowledge graphs constructed from nearly 900 papers. Domingo-Fernández et al. (Domingo-Fernández et al., 2021) create the COVID-19 Knowledge Graph, a comprehensive cause-and-effect network constructed from the scientific literature on the coronavirus. Additionally, Turki et al. (Turki et al., 2022) use knowledge graphs to assess and validate the portion of Wikidata related to COVID-19 epidemiology using an automatable task set. Pressat Laffouilhère et al. (Pressat Laffouilhère et al., 2022) develop OntoBioStat, a domain ontology related to covariate selection and bias in biostatistics, which can help interpret significant statistical associations between variables.
@@
Environmental Health. Fecho et al. (Fecho et al., 2021) develop ROBOKOP, a biomedical knowledge graph-based system, to validate associations between workplace chemical exposures and immune-mediated diseases. Wolffe et al. (Wolffe et al., 2020) propose using knowledge graphs in systematic evidence mapping in environmental health. This approach overcomes the limitations of rigid data tables by offering a more suitable model for handling the highly connected and complex nature of environmental health data.
@@
Health Policy and Management. Wu et al. (Wu, 2021) have analyzed the COVID-19 epidemic situation using a knowledge graph of patient activity. This method enables in-depth study of the transmission process, analysis of key nodes, and tracing of activity tracks. Meanwhile, Yu et al. (Yu et al., 2022) develop a chronic management system, which combines knowledge graphs and big data to optimize the management of chronic diseases in children. This system enhances treatment and resource utilization while conforming to the requirements of the Chronic Care Model.
@@
Social and Behavioral Health. There have been several interesting research studies. Cao et al. (Cao et al., 2020b) build a high-level suicide-oriented knowledge graph combined with deep neural networks for detecting suicidal ideation on social media platforms. Also, Liu et al. (Liu et al., 2020) conduct a bibliometric analysis of driver behavior research. Additionally, Wang et al. (Wang et al., 2022d) create an analysis framework for interpreting causal associations in emotional logic. They introduce a knowledge graph into appraisal theories, improving human emotional inference.
5. Promise and Outlook
The potential impact of comprehensive and fine-grained HKGs on biomedical research and clinical practice is significant. By integrating vast amounts of biomedical knowledge from multiple domains, HKGs can facilitate the discovery of new disease mechanisms and the identification of novel drug targets. They also help to enable personalized medicine by identifying patient subgroups with shared disease mechanisms. The recent success of large language models (a.k.a. foundation models) such as ChatGPT offers promising opportunities in capturing such semantics from the biomedical context (Agrawal et al., 2022; Singhal et al., 2022; Nath et al., 2022; Moor et al., 2023), enabling the construction of unprecedentedly comprehensive HKGs. In turn, HKGs also help improve LLMs by providing accurate and contextualized knowledge to regularize the generated content. This is particularly useful in evaluating LLMs in biomedical applications and addressing the problem of hallucination in critical areas.
6. Conclusion
@@
Healthcare knowledge graphs have emerged as a promising approach for capturing and organizing medical knowledge in a structured and interpretable way. This comprehensive review paper provides an overview of the current state of HKGs, including their construction, utilization models, and applications in healthcare. Furthermore, the paper discusses the potential future developments of HKGs. In conclusion, HKGs have played a significant role in advancing health research. With the advent of LLMs, there are even more opportunities to combine HKGs and LLMs to reduce the generation of false or unreliable content. We hope that our comprehensive review of this field offers a helpful perspective for future reference.

### v5

@@ 3.3 Clinical Decision Support
Electronic Health Record (EHR) contains essential patient information such as disease diagnoses, prescribed medications, and test results. However, the sparsity of EHR data largely restricts the ability of deep learning approaches. To overcome this drawback, KGs have been applied to incorporate prior medical knowledge for these deep learning models to better support the downstream prediction tasks.
3.3.1 Intermediate Steps to Advance Prediction Models
@@
ICD Coding aims to extract diagnosis and procedure codes from clinical notes which are often raw texts 142; 143; 144; 145. It is often challenging, as the size of the candidate target codes can be large and the distribution of the codes is often long-tailed 146. To overcome this, Xie et al. 147 and Cao et al. 148 propose to leverage KGs as distant supervision 149; 58, and inject the label information via structured KG propagation by leveraging graph convolution networks 150 to learn the correlations among medical codes. Besides, Lu et al. 151 propose to leverage KGs and the co-occurrence graph among clinical nodes simultaneously with a knowledge aggregation module to boost the ICD coding performance. Ren et al. 152 design a learning curriculum based on the hierarchical structure of the code to address the highly imbalanced label distribution issue and balance between frequent and rare labels. Overall, injecting additional knowledge with graph neural networks offers a way to mitigate the imbalanced label distribution issue and thus better.
@@
Entity and Relation Extraction helps convert the rich unstructured or semi-structured data in health records text into structured data that can be more easily processed, understood, and utilized by clinicians and algorithms. Specifically, entity extraction aims to identify entity mentions from clinical-free texts. There are two key steps for entity extraction, i.e., named entity recognition (NER) and disambiguation (NED). By leveraging additional KGs, Yuan et al. 153 inject additional knowledge from the KGs for entity linking and proposed post-pruning and Thresholding to improve the efficiency and remove the effect of unlinkable entity mention. Fries et al. 73 leverage clinical ontologies to provide weak supervision sources to create additional training data for clinical entity disambiguation. Besides, relation extraction aims to identify and classify relationships between entities in unstructured text, which facilitates understanding complex biological processes, drug interactions, and disease mechanisms. To incorporate the external KG, several works 154; 155 proposed additional post-training steps to align the language models with biomedical knowledge. Hong et al. 156 construct embeddings for a wide range of codified concepts from EHRs to identify relevant features related to a disease of interest, and Lin et al. 157 design a co-training scheme to jointly learn from text and KGs for extracting and classifying disease-disease relations. In summary, fusing KGs with language models can flexibly accommodate missing data types and bring additional performance gains, especially for those rare entities and relations.
3.3.2 Evidence Generation for Risk Prediction Models
@@
Disease Prediction aims to predict the potential diseases of a given patient with his past clinical records. To assist the diagnosis with additional knowledge, GRAM 158 and KAME 159 utilize a medical ontology 160 where the leaf nodes are the medical codes found in EHR data, and their ancestors are more general categories. By incorporating information from medical ontologies into deep learning models via neural attention, these approaches learn better embeddings for different medical concepts to alleviate the data scarcity bottleneck. 161; 162 further consider the domain-specific KG KnowLife 163 to enrich the embeddings of medical entities with their neighbors on the KG. These approaches mainly directly update the embeddings of different concepts to improve the feature learning, but may be at the risk of ignoring the high-level order information from the KG. To tackle this drawback, Ye et al. 164 explicitly exploit paths in KG from the observed symptoms to the target disease to model the personalized information for diverse patients. Xu et al. 165 design a self-supervised learning approach to pre-train a graph attention network for learning the embedding of medical concepts and completing the KG simultaneously. These approaches better harness the structure information, and often lead to better performance than the pure embedding-based knowledge integration techniques.
@@
Treatment Recommendation aims to recommend personalized medications to patients based on their individual health conditions, which can help physicians select the most effective medications for their patients, and improve treatment outcomes 166; 167; 168. To effectively exploit external knowledge, Shang et al. 169 use drug ontologies to design additional pretraining loss and directly improve the representation of drugs, and several studies 170; 171 attempt to extract the additional drug interaction graphs to model the negative side effects of specific drug pairs and reduce the possibility of recommending negative drug-drug interaction combinations. Besides, Wu et al. 172 leveraged ontologies to improve the drug representations, and facilitates drug recommendation under a more challenging few-shot setting.
3.4 Public Health
@@
Public Health research can significantly benefit from HKGs. KGs can help organize, structure, and formalize extensive information from diverse and heterogeneous sources. This allows researchers to analyze data, reason about factors, and make decisions on a larger scale.
@@
Epidemiology. The field of epidemiology has seen an increased use of KGs to analyze and understand the spread of diseases. A study by Gao et al. 173 analyzes the research and development trends of wastewater-based epidemiology (WBE) using KGs constructed from nearly 900 papers. Domingo-Fernández et al. 174 create the COVID-19 KG, a comprehensive cause-and-effect network constructed from the scientific literature on the coronavirus. Additionally, Turki et al. 175 use KGs to assess and validate the portion of Wikidata related to COVID-19 epidemiology using an automatable task set. Pressat Laffouilhère et al. 176 develop OntoBioStat, a domain ontology related to covariate selection and bias in biostatistics, which can help interpret significant statistical associations between variables.
@@
Environmental Health. Fecho et al. 177 develop ROBOKOP, a biomedical KG-based system, to validate associations between workplace chemical exposures and immune-mediated diseases. Wolffe et al. 178 propose using KGs in systematic evidence mapping in environmental health. This approach overcomes the limitations of rigid data tables by offering a more suitable model for handling the highly connected and complex nature of environmental health data.
@@
Health Policy and Management. Wu et al. 179 have analyzed the COVID-19 epidemic situation using a KG of patient activity. This method enables in-depth study of the transmission process, analysis of key nodes, and tracing of activity tracks. Meanwhile, Yu et al. 180 develop a chronic management system, which combines KGs and big data to optimize the management of chronic diseases in children. This system enhances treatment and resource utilization while conforming to the requirements of the Chronic Care Model.
@@
Social and Behavioral Health. Cao et al. 181 build a high-level suicide-oriented KG combined with deep neural networks for detecting suicidal ideation on social media platforms. Also, Liu et al. 182 conduct a bibliometric analysis of driver behavior research. Additionally, Wang et al. 183 create an analysis framework for interpreting causal associations in emotional logic. They introduce a KG into appraisal theories, improving human emotional inference.
Table 2: Main challenges in different types of problems.
Types of Problems Main Challenges
Basic Science Research - Limited integration of KGs with LLMs for structural and functional predictions of molecules and genes 184.
- Difficulty in representing high-dimensional data (e.g., protein structures) in KGs 185.
- Lack of domain-specific fine-tuning of LLMs for bioinformatics tasks 186.
Pharmaceutical Development - Incomplete or inconsistent drug-related knowledge representation in KGs 187.
- Challenges in combining structured KG data with unstructured LLM outputs (e.g., literature mining) for assisting drug discovery 92.
Clinical Decision Support - Difficulty in ensuring explainability and reliability of predictions 188.
- Lack of harmonization between real-time clinical data and static KGs for dynamic decision-making 32.
Public Health - Limited ability to model causal relationships and complex interactions in public health domains using KGs 189.
- Difficulty in integrating population-level data with KGs for real-time surveillance. 190
4 Challenge, Promise, and Outlook
Current Challenges for Adopting KGs to Biomedical Research. Injecting KGs into clinical problems faces significant challenges across domains. In basic science research, the integration of KGs and LLMs for molecular predictions is limited by difficulties in representing high-dimensional data like protein structures and a lack of domain-specific fine-tuning 184; 185; 186. In pharmaceutical development, inconsistent drug-related knowledge representation and challenges in combining KGs with unstructured LLM outputs hinder drug discovery 187; 92. Clinical decision support struggles with ensuring explainability, reliability, and harmonizing real-time clinical data with static KGs for adaptive decision-making 188; 32. Public health faces limitations in modeling causal relationships and integrating population-level data with KGs for real-time surveillance due to scalability and standardization issues 189; 190. Addressing these barriers is vital for leveraging KGs to advance clinical and public health outcomes.
Future Directions The potential impact of comprehensive and fine-grained HKGs on biomedical research and clinical practice is significant. By integrating vast amounts of biomedical knowledge from multiple domains, HKGs can facilitate the discovery of new disease mechanisms and the identification of novel drug targets. They also help to enable personalized medicine by identifying patient subgroups with shared disease mechanisms. The recent success of LLMs such as ChatGPT offers promising opportunities in capturing such semantics from the biomedical context 72; 191; 192; 193, enabling the construction of unprecedentedly comprehensive HKGs. In turn, HKGs also help improve LLMs by providing accurate and contextualized knowledge to regularize the generated content. This is particularly useful in evaluating LLMs in biomedical applications and addressing the problem of hallucination in critical areas. For developers of novel biomedical informatics methods, this symbiotic relationship between HKGs and LLMs presents both opportunities and challenges. One key area for innovation is the design of frameworks that allow real-time updating of HKGs with the latest clinical and research data, enabling dynamic decision-making. Developers must also address the computational and representational challenges of integrating high-dimensional data, such as genomic sequences and protein structures, into HKGs while ensuring semantic consistency. Moreover, constructing HKGs that incorporate causal reasoning capabilities could significantly enhance their utility in understanding complex disease mechanisms and predicting therapeutic outcomes. Another promising avenue lies in leveraging fine-tuned LLMs to extract domain-specific insights from unstructured biomedical literature and seamlessly integrate them into HKGs. This process, however, requires rigorous validation pipelines to ensure the accuracy and relevance of the extracted information. Developers should also explore strategies for aligning LLM outputs with ontologies and existing knowledge schemas to maximize interoperability and reuse. Finally, as HKGs are increasingly utilized to support clinical decision-making, developers must prioritize explainability and fairness in their design. This includes developing visualization tools and interactive interfaces that allow clinicians and researchers to interpret the outputs of HKG-powered systems effectively. Similarly, integrating multilingual capabilities and context-aware reasoning into HKGs will be essential for addressing global healthcare challenges and ensuring equitable access to cutting-edge biomedical insights.
5 Conclusion
@@
HKGs have emerged as a promising approach for capturing and organizing medical knowledge in a structured and interpretable way. This comprehensive review paper provides an overview of the current state of HKGs, including their construction, modeling, and applications in healthcare. Furthermore, the paper discusses potential future developments of HKGs. In conclusion, HKGs have played a significant role in advancing health research. With the advent of LLMs, there are even more opportunities to combine HKGs and LLMs to reduce the generation of false or unreliable content. We hope that our comprehensive review of this field offers a helpful perspective for future reference.
Acknowledgements
This research was partially supported by the National Science Foundation under Award Number 2319449 and Award Number 2312502, as well as the National Institute Of Diabetes And Digestive And Kidney Diseases of the National Institutes of Health under Award Number K25DK135913. Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and do not necessarily represent the views, either expressed or implied, of the National Science Foundation, National Institutes of Health, or the U.S. government. The authors wish to thank the editors and reviewers for their valuable efforts and suggestions.